SpaceRef

SpaceRef


ESA Venus Express Status Report No. 90

Status Report From: ESA Venus Express Mission
Posted: Wednesday, October 17, 2007

image

06 Aug 2007 12:00 Report for Period 29 July to 04 August 2007

Routine observations took place during this reporting period, occurring as planned. On DOY 211 ASPERA suffered a failure to boot-up, and a complete science observation failed as a result.


MET (Day) Date DOY VPER# Main Activity
628 29/07/07 210 464 Routine Operations
629 30/07/07 211 465

Routine Operations
ASPERA suffered a hang-up and failed to boot-up, one science observation lost

630 31/07/07 212 466

Routine Operations
STR A power-cycled to clear candidate hot pixel list (AR-72).

631 01/08/07 213 467

Routine Operations

632 02/08/07 214 468

Routine Operations
TM bit rate change to 182.8 Kbps before AOS. BSR pass on DSN before CEB.

633 03/08/07 215 469

Routine Operations
BSR pass on DSN before CEB

634 04/08/07 216 470 Routine Operations

At the end of the last Cebreros pass in the reporting period (DOY 216, 18:00z) Venus Express was orbiting Venus at 47.1 million km from the Earth. The one-way signal travel time was 156 seconds.

Payload Activities

ASPERA
On DOY 211, the instrument failed to boot up as commanded, and appeared to remain in Safe Mode. This caused all subsequent TCs related to the science run not to be executed, and thus the observation to be lost.

MAG
The instrument was regularly operated as part of the routine plan.

PFS
Following the completion of the planned tests last week, there has not been any further operation of PFS.

SPICAV
The instrument was regularly operated as part of the routine plan.

VeRA
Two Bi-Static Radar passes were scheduled during this period, on DOY 214 and 215. These BSR activities were the last ones for the year, all the others in the schedule having been cancelled due to the S-band attenuation problem.

VIRTIS
The instrument was regularly operated as part of the routine plan.

VMC
The instrument is operated as part of the routine plan.

Future Milestones

  • VIRTIS Movie #2 scheduled in the second week of August.
  • Low PLL bandwidth acquisition tests with LGA and S-band over NNO scheduled for mid-August.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks